6-Estimated data for Lorien Company for Year 1 are as follows:

Direct materials to be purchased...................................................... $35,000

Indirect labor.......................................................................................... 10,000

Wages of janitors in corporate office building................................. 47,000

Factory building rent............................................................................. 60,000

Depreciation on production equipment............................................. 20,000

Administrative office supplies............................................................. 25,000

Estimated direct labor hours: 40,000 hours

Actual data for Year 1 are as follows:

Total manufacturing overhead: $110,000

Direct labor hours: 35,000 hours

The predetermined manufacturing overhead rate is determined on the basis of direct labor hours. Which ONE of the following statements is TRUE?

Applied overhead was MORE than actual overhead by $12,250

Applied overhead was LESS than actual overhead by $20,000

Applied overhead was LESS than actual overhead by $10,000

Applied overhead was MORE than actual overhead by $20,000

Applied overhead was LESS than actual overhead by $12,250

Applied overhead was MORE than actual overhead by $31,250

Applied overhead was LESS than actual overhead by $31,250

Applied overhead was MORE than actual overhead by $10,000
Answer #1

Estimated overhead = 10000+60000+20000 = 90000

Predetermine overhead rate = 90000/40000 = 2.25 per labour hour

Applied overhead = 35000*2.25 = 78750

Actual overhead = 110000

Difference = 110000-78750 = 31250

So answer is Applied overhead was LESS than actual overhead by $31,250